Please help.
What is the point-slope form of an equation with the slope 3/5 that passes through the point (10,-2) ?

Answer by stanbon(48516)

What is the point-slope form of an equation with the slope 3/5 that passes through the point (10,-2) ?
The equation of the line will be y=mx+b where y=-2, x=10, m=(3/5)
Thus -2=(3/5)(10)+b
-2=6+b
b=-8
EQUATION: y=(3/5)x-8
Cheers,
Stan H.